there are 2 male students to every 7 female students at a local school. let x be the number of female students when there are 27 male students. selecet the proportion that can be used to find x.
Found 2 solutions by shree840, stanbon:
Answer by shree840(260)

male/female=2/7=27/x
x=7*27/2=189/2=94.5 ANS (I will take the half girl)

Answer by stanbon(75887)

there are 2 male students to every 7 female students at a local school. let x be the number of female students when there are 27 male students. selecet the proportion that can be used to find x.
----------------------
female/male = 7/2
---
7:2 is the same as 7k to 2k
---
If # of females is 7k = x, then k = x/7; and # of males is 2k = 2(x/7)
----
Equation:
# of males = # of males
2x/7 = 27/1
---
2x = 7*27
--
x = (7*27)/2 = 94.5 is the # of female students
===========
Cheers,
Stan H.